# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 7.2. Banach Roe-type operator algebras
In this subsection we associate to any uniformly locally finite coarse structure $(X,\operatorname{\mathbb{E}})$ a family of Banach algebras, that cover Roe-type Banach algebras with compact coefficients considered by Špakula and Willet [SW17], as well as $\ell^{p}$ Roe-type algebras of Chung and Li [CL18]. We prove that they can be naturally viewed as the reduced Banach algebra of a principal étale groupoid, so parts of our theory apply.
Let $\operatorname{\mathbb{E}}$ be a (unital) coarse structure on a set $X$ in the sense of [STY02], i.e. $\operatorname{\mathbb{E}}$ is a collection of subsets of the pair groupoid $X\times X$ which contains all finite subsets, and is closed under inverses, products, unions and formation of subsets. Elements of $\operatorname{\mathbb{E}}$ are called entourages. The coarse space $(X,\operatorname{\mathbb{E}})$ is called unital if $\operatorname{\mathbb{E}}$ contains the diagonal $\{(x,x):x\in X\}$. We will assume $(X,\operatorname{\mathbb{E}})$ is uniformly locally finite, which means that for any $E\in\operatorname{\mathbb{E}}$ the range and domain of $E$ is finite, i.e. the following number
@@
is finite. As explained in [STY02, page 810], up to coarse equivalence coarse structures coming from metric spaces are exactly countably generated unital coarse structures. Then uniform local finiteness corresponds to what is called bounded geometry for metric spaces. In this sense, our setting covers those in [SW17] or [CL18].
Let $\mathcal{E}$ be a Banach space with a fixed normalised (not necessarily countable) Schauder basis $\{e_{i}\}_{i\in I}$. For every finite $I_{0}\subseteq I$ we denote by $M_{I_{0}}(\mathcal{E})\cong M_{|I_{0}|}(\mathbb{F})$ the set of operators on the finite dimensional space $\operatorname{span}\{e_{i}:i\in I_{0}\}$. We view $M_{I_{0}}(\mathcal{E})$ as a subaglebra of $\operatorname{\mathbb{B}}(\mathcal{E})$ and we put $M_{I}(\mathcal{E}):=\bigcup_{\stackrel{{\scriptstyle I_{0}\subseteq I}}{{\text{finite}}}}M_{I_{0}}\subseteq\operatorname{\mathbb{B}}(\mathcal{E})$. Up to an equivalence of the norm in $\mathcal{E}$, we may and we will assume that the unit $1_{I_{0}}$ in $M_{I_{0}}(\mathcal{E})$ has norm one as an element of $\operatorname{\mathbb{B}}(\mathcal{E})$, cf. [Daw06]. We will refer to such a Schauder basis as standardized.
@@
The collection of all such controlled propagation matrices is denoted $\mathbb{F}[X;\mathcal{E}]$.
@@
Fix a finite $I_{0}\subseteq I$ and an entourage $E\in\operatorname{\mathbb{E}}$. Bounded maps from $X$ to $M_{I_{0}}(\mathcal{E})$ supported on $r(E):=\{x:(x,y)\in E\}$, i.e. elements of $\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$, when identified with diagonal matrices, are elements of $\mathbb{F}[X;\mathcal{E}]$, referred to as multiplication operators. If in addition the entourage $E$ is a bisection in $X\times X$, the characteristic function of $E$ multiplied by the unit $1_{I_{0}}$ in $M_{I_{0}}(\mathcal{E})$ is an element of $\mathbb{F}[X;\mathcal{E}]$, referred to as a partial translation. Namely, for a bisection entourage $E$ and any $f\in\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$ the formulas
@@
Let $T\in\mathbb{F}[X;\mathcal{E}]$ and put $N:=n(\operatorname{supp}(T))^{2}+1<\infty$. Then there are pairwise bisection entourages $E_{1},\ldots,E_{N}\subseteq\operatorname{supp}(T)$ and functions $f_{1},\ldots,f_{N}\in\ell^{\infty}(X,M_{I_{0}}(\mathcal{E}))$, for some finite $I_{0}\subseteq I$, with $\operatorname{supp}_{x\in X,k=1,\ldots,N}\|f_{k}(x)\|\leq\sup_{x,y}\|T_{x,y}\|$ and such that
@@
Let $I_{0}\subseteq I$ be a finite set such that $T_{x,y}\subseteq M_{I_{0}}(\mathcal{E})$ all $x,y\in X$. By [Roe03, Lemma 4.10], see also [STY02, Lemma 2.7], there are pairwise bisection entourages $E_{1},\ldots,E_{N}$ such that $\operatorname{supp}(T)=\bigsqcup_{i=1}^{N}E_{i}$. For each $i=1,\ldots,n$ define $f_{i}\in\ell^{\infty}(r(E_{i}),M_{I_{0}}(\mathcal{E}))$ by $f_{i}(x):=T_{x,y}$ if $(x,y)\in E$ for some $y\in X$ and $f_{i}(x)=0$ otherwise. Then the displayed formula holds. ∎
@@
The usual matrix operations and the algebra structure of $\operatorname{\mathbb{B}}(\mathcal{E})$ make $\mathbb{F}[X;\mathcal{E}]$ into an algebra. Moreover, for every $p\in[1,+\infty]$ a matrix $T\in\mathbb{F}[X;\mathcal{E}]$ defines via
$(T_{p}\xi)(x):=\sum_{y\in X}T_{x,y}\xi(y),\qquad\xi\in\ell^{p}(X,\mathcal{E}),$
a bounded operator $T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ with $\|T_{p}\|\leq\sup_{x,y}\|T_{x,y}\|\cdot n(\operatorname{supp}(T))^{2}+1$. The assignment $\mathbb{F}[X;\mathcal{E}]\ni T\mapsto T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is an injective algebra homomorphism.
@@
Let $E\in\operatorname{\mathbb{E}}$ be a bisection entourage and let $f\in\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$, for a finite $I_{0}\subseteq I$. The norm of $(fE^{I_{0}})_{p}$, as an operator on $\ell^{p}(X,\mathcal{E})$, is not greater than $\sup_{x\in X}\|f(x)\|$ due to our standing normalization assumption that $1_{I_{0}}\in\mathcal{B}(\mathcal{E})$ has norm one (which implies that $(E^{I_{0}})_{p}$ has norm one). Thus applying Lemma 7.18 to $T\in\mathbb{F}[X;\mathcal{E}]$ we get that $T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is bounded with the prescribed estimate on the norm.
It is clear that the map $\mathbb{F}[X;\mathcal{E}]\ni T\mapsto T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is linear and injective. To see it is multiplicative, in view of Lemma 7.18, it suffices to consider only two elements of the form $f_{1}E_{1}^{I_{1}},f_{2}E_{2}^{I_{2}}\in\mathbb{F}[X;\mathcal{E}]$ where $f_{1}\in\ell^{\infty}(X,M_{I_{1}}(\mathcal{E}))$, $f_{2}\in\ell^{\infty}(X,M_{I_{2}}(\mathcal{E}))$, $I_{1},I_{2}\subseteq I$ are finite, and $E_{1},E_{2}\in\operatorname{\mathbb{E}}$ are bisections. Then we have
(7.14) $(f_{1}E_{1}^{I_{1}})(f_{2}E_{2}^{I_{2}})=f_{1}(f_{2}\circ h_{E_{1}})(E_{1}E_{2})^{I_{1}\cap I_{2}}$
@@
For any non-empty $P\subseteq[1,\infty]$ we define the $L^{P}$$-\mathcal{E}$ Roe-type Banach algebra of $(X,\operatorname{\mathbb{E}})$ as the completion $F^{P}[X;\mathcal{E}]=\overline{\mathbb{F}[X;\mathcal{E}]}^{\|\cdot\|_{L^{P},\mathcal{E}}}$ in the norm $\|T\|_{L^{P},\mathcal{E}}:=\sup_{p\in P}\|T_{p}\|$. When $P=\{p\}$ is a singleton we omit writing braces in the above notation.
The translation groupoid associated to the coarse space $(X,\operatorname{\mathbb{E}})$ is defined as the union of closures of entourages in the Stone–Čech compactifaction of the discrete space $X\times X$, see [STY02, 3.2], [Roe03, Definition 10.19]. That is as a topological space we put
@@
under which for any non-empty $P\subseteq[1,\infty)$ (or $P\subseteq[1,\infty]$ when $\mathcal{E}$ is reflexive), the Roe-type algebra $F^{P}[X;\mathcal{E}]$ becomes a reduced $S$-graded groupoid Banach algebra of $\mathcal{G}$ where $S$ consists of $\overline{E}\times\{(i,j)\}$, $E\in S^{X}$, $i,j\in I$, and the unit space $\beta(X)$. In particular, the diagonal $C_{0}(\beta(X)\times I)$ is a maximal abelian subalgebra, $C_{0}(\beta(X)\times I)\subseteq F^{P}[X;\mathcal{E}]$ has the intersection property, and $F^{P}[X;\mathcal{E}]$ is simple if and only if the coarse space $(X,\operatorname{\mathbb{E}})$ is simple.
@@
where $f(x)_{i,j}:=\overline{f}(x,y,i,j)$ if $(x,y)\in E$ for some $y\in X$, and $f(x)=0$ otherwise. Using this and Lemma 7.18 one sees that the maps
@@
integrate to a linear isomorphism $C_{c}(\mathcal{G})\cong\mathbb{F}[X;\mathcal{E}]$. In view of (7.14) it readily follows that this isomorphism is multiplicative.
Let $p\in[1,\infty]$. Composing $C_{c}(\mathcal{G})\cong\mathbb{F}[X;\mathcal{E}]$ with the monomorphism from Proposition 7.19 we get a monomorphism $\Lambda_{p}:C_{c}(\mathcal{G})\to\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ such that
$[\Lambda_{p}(f)\xi](x)=\sum_{y\in X}f(x,y,i,j)e_{i}\otimes e_{j}^{*}(\xi(y)),\qquad f\in C_{c}(\mathcal{G}),\xi\in\ell^{p}(X,\mathcal{E}),$
where $e_{i}\otimes e_{j}^{*}\in\mathcal{B}(\mathcal{E})$ is the contractive one-dimensional projection given by $(e_{i}\otimes e_{j}^{*})(\sum_{k\in I}\xi_{k}e_{k})=\xi_{j}e_{i}$. For $f$ supported on $\overline{E}\times\{(i,j)\}\in S$ we clearly have $\|\Lambda_{p}(f)\|=\|f\|_{\infty}$, and so $\Lambda_{p}$ extends to a representation $\Lambda_{p}:F^{S}(\mathcal{G})\to F^{p}[X;\mathcal{E}]$. For $(y,j)\in X\times I$ we let $1_{(y,j)}\in\ell^{p}(X,\mathcal{E})$ be such that $1_{(y,j)}(y):=e_{j}$ and $01_{(y,j)}(x):=$ 0 for $x\in X\setminus\{y\}$. Then putting
@@
we get a contractive linear map $j$ from $F^{P}[X;\mathcal{E}]$ to $C_{b}(\mathcal{G}_{0})$, equipped with the norm $\|\cdot\|_{\infty}$. For $f\in C_{c}(\mathcal{G})$ we have $j(\Lambda_{p}(f))=f|_{\mathcal{G}_{0}}$ which has a unique continuous extension to $\mathcal{G}$ (which is $f$) and therefore it may be treated as an element of $C_{c}(\mathcal{G})$. Thus we get that on a dense subalgebra $j$ takes values in $C_{c}(\mathcal{G})$ and therefore we may assume that $j$ maps $F^{p}[X;\mathcal{E}]$ to $C_{0}(\mathcal{G})$. In particular, $F^{p}[X;\mathcal{E}]$ is a groupoid Banach algebra. If $p<\infty$, then $\{1_{(y,j)}\}_{(y,j)\in X\times I}$ is a Schauder basis for $\ell^{p}(X,\mathcal{E})$, and then (7.15) implies that the map $j$ is injective on $F^{p}[X;\mathcal{E}]$ and hence $F^{P}[X;\mathcal{E}]$ is reduced. For $p=\infty$ this is not so clear. However, when $\mathcal{E}$ is reflexive, then $\{e_{i}^{*}\}_{i\in I}$ is a Schauder basis for the dual space $\mathcal{E}^{*}$, and the involution on $C_{c}(\mathcal{G})$ induces an isometric anti-linear isomorphism $F^{\infty}[X;\mathcal{E}]\cong F^{1}[X;\mathcal{E}^{*}]$. Thus then $F^{\infty}[X;\mathcal{E}]$ is reduced because we have just proved that $F^{1}[X;\mathcal{E}^{*}]$ is. The assertion for the set $P\subseteq[1,\infty]$ follows now from Lemma 3.12.
@@
Using that $C_{0}(\beta(X)\times I)\subseteq F^{P}[X;\mathcal{E}]$ has the intersection property one can show that ideals in $F^{P}[X;\mathcal{E}]$ in which controlled propagation operators are dense, are exactly those generated by open $\mathcal{G}$-invariant subsets of $\beta(X)\times I$ (which in turn correspond to ideals in the coarse structure $(X,\operatorname{\mathbb{E}})$). In this way one could generalize [CWT04, Theorems 6.3, 6.4] from uniform Roe $C^{*}$-algebras to algebras of the form $F^{P}[X;\mathcal{E}]$.

### v3

@@ 7.2. Banach Roe-type operator algebras
In this subsection, we associate to any uniformly locally finite coarse structure $(X,\operatorname{\mathbb{E}})$ a family of Banach algebras that cover Roe-type Banach algebras with compact coefficients considered by Špakula and Willett [SW17], as well as $\ell^{p}$ Roe-type algebras of Chung and Li [CL18]. We prove that they can be naturally viewed as a reduced Banach algebra of a principal étale groupoid, so parts of our theory apply.
Let $\operatorname{\mathbb{E}}$ be a (unital) coarse structure on a set $X$ in the sense of [STY02], that is, $\operatorname{\mathbb{E}}$ is a collection of subsets of the pair groupoid $X\times X$ which contains all finite subsets, and is closed under inverses, products, unions and formation of subsets. Elements of $\operatorname{\mathbb{E}}$ are called entourages. The coarse space $(X,\operatorname{\mathbb{E}})$ is called unital if $\operatorname{\mathbb{E}}$ contains the diagonal $\{(x,x):x\in X\}$. We will assume that $(X,\operatorname{\mathbb{E}})$ is uniformly locally finite, which means that for any $E\in\operatorname{\mathbb{E}}$ the range and domain of $E$ is finite, that is, the following number
@@
is finite. As explained in [STY02, page 810], up to coarse equivalence, coarse structures coming from metric spaces are exactly countably generated unital coarse structures. In this case uniform local finiteness corresponds to what is called bounded geometry for metric spaces. In this sense, our setting covers those in [SW17] or [CL18].
Let $\mathcal{E}$ be a Banach space with a fixed normalized (not necessarily countable) Schauder basis $\{e_{i}\}_{i\in I}$. For every finite $I_{0}\subseteq I$, we denote by $M_{I_{0}}(\mathcal{E})\cong M_{|I_{0}|}(\mathbb{F})$ the set of operators on the finite dimensional space $\operatorname{span}\{e_{i}:i\in I_{0}\}$. We view $M_{I_{0}}(\mathcal{E})$ as a subaglebra of $\operatorname{\mathbb{B}}(\mathcal{E})$ and we put $M_{I}(\mathcal{E}):=\bigcup_{\stackrel{{\scriptstyle I_{0}\subseteq I}}{{\text{finite}}}}M_{I_{0}}\subseteq\operatorname{\mathbb{B}}(\mathcal{E})$. Up to an equivalence of the norm in $\mathcal{E}$, we may and shall assume that the unit $1_{I_{0}}$ in $M_{I_{0}}(\mathcal{E})$ has norm one as an element of $\operatorname{\mathbb{B}}(\mathcal{E})$, cf. [Daw06]. We will refer to such a Schauder basis as standardized.
@@
The collection of all such controlled propagation matrices is denoted by $\mathbb{F}[X;\mathcal{E}]$.
@@
Fix a finite $I_{0}\subseteq I$ and an entourage $E\in\operatorname{\mathbb{E}}$. Bounded maps from $X$ to $M_{I_{0}}(\mathcal{E})$ supported on $r(E):=\{x:(x,y)\in E\}$, that is elements of $\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$, when identified with diagonal matrices, are elements of $\mathbb{F}[X;\mathcal{E}]$, referred to as multiplication operators. If in addition the entourage $E$ is a bisection in $X\times X$, the characteristic function of $E$ multiplied by the unit $1_{I_{0}}$ in $M_{I_{0}}(\mathcal{E})$ is an element of $\mathbb{F}[X;\mathcal{E}]$, referred to as a partial translation. Namely, for a bisection entourage $E$ and any $f\in\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$, the formulas
@@
Let $T\in\mathbb{F}[X;\mathcal{E}]$ and put $N:=n(\operatorname{supp}(T))^{2}+1<\infty$. Then there are pairwise disjoint bisection entourages $E_{1},\ldots,E_{N}\subseteq\operatorname{supp}(T)$ and functions $f_{1},\ldots,f_{N}\in\ell^{\infty}(X,M_{I_{0}}(\mathcal{E}))$, for some finite $I_{0}\subseteq I$, with $\operatorname{supp}_{x\in X,k=1,\ldots,N}\|f_{k}(x)\|\leq\sup_{x,y}\|T_{x,y}\|$ and such that
@@
Let $I_{0}\subseteq I$ be a finite set such that $T_{x,y}\subseteq M_{I_{0}}(\mathcal{E})$ all $x,y\in X$. By [Roe03, Lemma 4.10], see also [STY02, Lemma 2.7], there are pairwise disjoint bisection entourages $E_{1},\ldots,E_{N}$ such that $\operatorname{supp}(T)=\bigsqcup_{i=1}^{N}E_{i}$. For each $i=1,\ldots,n$ define $f_{i}\in\ell^{\infty}(r(E_{i}),M_{I_{0}}(\mathcal{E}))$ by $f_{i}(x):=T_{x,y}$ if $(x,y)\in E$ for some $y\in X$ and $f_{i}(x)=0$ otherwise. Then the displayed formula holds. ∎
@@
The usual matrix operations and the algebra structure of $\operatorname{\mathbb{B}}(\mathcal{E})$ turn $\mathbb{F}[X;\mathcal{E}]$ into an algebra. Moreover, for every $p\in[1,+\infty]$, a matrix $T\in\mathbb{F}[X;\mathcal{E}]$ defines a bounded operator $T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ via
$(T_{p}\xi)(x):=\sum_{y\in X}T_{x,y}\xi(y),\qquad\xi\in\ell^{p}(X,\mathcal{E}).$
Then $\|T_{p}\|\leq\sup_{x,y}\|T_{x,y}\|\cdot n(\operatorname{supp}(T))^{2}+1$. The assignment $\mathbb{F}[X;\mathcal{E}]\ni T\mapsto T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is an injective algebra homomorphism.
@@
Let $E\in\operatorname{\mathbb{E}}$ be a bisection entourage and let $f\in\ell^{\infty}(r(E),M_{I_{0}}(\mathcal{E}))$, for a finite $I_{0}\subseteq I$. The norm of $(fE^{I_{0}})_{p}$, as an operator on $\ell^{p}(X,\mathcal{E})$, is not greater than $\sup_{x\in X}\|f(x)\|$ due to our standing normalization assumption that $1_{I_{0}}\in\mathcal{B}(\mathcal{E})$ has norm one, which implies that $(E^{I_{0}})_{p}$ has norm one. Thus, applying Lemma 7.18 to $T\in\mathbb{F}[X;\mathcal{E}]$, we get that $T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is bounded with the prescribed estimate on the norm.
It is clear that the map $\mathbb{F}[X;\mathcal{E}]\ni T\mapsto T_{p}\in\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ is linear and injective. To see that it is multiplicative, in view of Lemma 7.18, it suffices to consider only two elements of the form $f_{1}E_{1}^{I_{1}},f_{2}E_{2}^{I_{2}}\in\mathbb{F}[X;\mathcal{E}]$ where $f_{1}\in\ell^{\infty}(X,M_{I_{1}}(\mathcal{E}))$, $f_{2}\in\ell^{\infty}(X,M_{I_{2}}(\mathcal{E}))$, $I_{1},I_{2}\subseteq I$ are finite, and $E_{1},E_{2}\in\operatorname{\mathbb{E}}$ are bisections. Then we have
(7.13) $(f_{1}E_{1}^{I_{1}})(f_{2}E_{2}^{I_{2}})=f_{1}(f_{2}\circ h_{E_{1}})(E_{1}E_{2})^{I_{1}\cap I_{2}},$
@@
For any non-empty $P\subseteq[1,\infty]$ we define the $L^{P}$-$\mathcal{E}$ Roe-type Banach algebra of $(X,\operatorname{\mathbb{E}})$ as the completion $F^{P}[X;\mathcal{E}]=\overline{\mathbb{F}[X;\mathcal{E}]}^{\|\cdot\|_{L^{P},\mathcal{E}}}$ in the norm $\|T\|_{L^{P},\mathcal{E}}:=\sup_{p\in P}\|T_{p}\|$. When $P=\{p\}$ is a singleton, we omit writing braces in the above notation.
The translation groupoid associated to the coarse space $(X,\operatorname{\mathbb{E}})$ is defined as the union of closures of entourages in the Stone–Čech compactifaction of the discrete space $X\times X$, see [STY02, 3.2], [Roe03, Definition 10.19]. That is, as a topological space, we put
@@
under which for any non-empty $P\subseteq[1,\infty)$ (or $P\subseteq[1,\infty]$ when $\mathcal{E}$ is reflexive), the Roe-type algebra $F^{P}[X;\mathcal{E}]$ becomes a reduced $S$-graded groupoid Banach algebra of $\mathcal{G}$ where $S$ consists of $\overline{E}\times\{(i,j)\}$, $E\in S^{X}$, $i,j\in I$, and the unit space $\beta(X)$.
In particular, the diagonal $C_{0}(\beta(X)\times I)$ is a maximal abelian subalgebra, $C_{0}(\beta(X)\times I)\subseteq F^{P}[X;\mathcal{E}]$ has the intersection property, and $F^{P}[X;\mathcal{E}]$ is simple if and only if the coarse space $(X,\operatorname{\mathbb{E}})$ is simple.
@@
where $f(x)_{i,j}:=\overline{f}(x,y,i,j)$ if $(x,y)\in E$ for some $y\in X$, and $f(x)=0$ otherwise. Using this and Lemma 7.18, one sees that the maps
@@
integrate to a linear isomorphism $C_{c}(\mathcal{G})\cong\mathbb{F}[X;\mathcal{E}]$. In view of (7.13), it readily follows that this isomorphism is multiplicative.
Let $p\in[1,\infty]$. Composing $C_{c}(\mathcal{G})\cong\mathbb{F}[X;\mathcal{E}]$ with the monomorphism from Proposition 7.19, we get a monomorphism $\Lambda_{p}:C_{c}(\mathcal{G})\to\operatorname{\mathbb{B}}(\ell^{p}(X,\mathcal{E}))$ such that
$\Lambda_{p}(f)\xi](x)=\sum_{y\in X}f(x,y,i,j)e_{i}\otimes e_{j}^{*}(\xi(y)),\qquad f\in C_{c}(\mathcal{G}),\xi\in\ell^{p}(X,\mathcal{E}),$
where $e_{i}\otimes e_{j}^{*}\in\mathcal{B}(\mathcal{E})$ is the contractive one-dimensional projection given by $(e_{i}\otimes e_{j}^{*})(\sum_{k\in I}\xi_{k}e_{k})=\xi_{j}e_{i}$. For $f$ supported on $\overline{E}\times\{(i,j)\}\in S$, we clearly have $\|\Lambda_{p}(f)\|=\|f\|_{\infty}$, and so $\Lambda_{p}$ extends to a representation $\Lambda_{p}:F^{S}(\mathcal{G})\to F^{p}[X;\mathcal{E}]$. For $(y,j)\in X\times I$, we let $1_{(y,j)}\in\ell^{p}(X,\mathcal{E})$ be such that $1_{(y,j)}(y):=e_{j}$ and $1_{(y,j)}(x):=0$ for $x\in X\setminus\{y\}$. Then putting
@@
we get a contractive linear map $j$ from $F^{P}[X;\mathcal{E}]$ to $C_{b}(\mathcal{G}_{0})$, equipped with the norm $\|\cdot\|_{\infty}$. For $f\in C_{c}(\mathcal{G})$, we have $j(\Lambda_{p}(f))=f|_{\mathcal{G}_{0}}$ which has a unique continuous extension to $\mathcal{G}$ (which is $f$) and therefore it may be treated as an element of $C_{c}(\mathcal{G})$. Thus we get that on a dense subalgebra $j$ takes values in $C_{c}(\mathcal{G})$ and therefore we may assume that $j$ maps $F^{p}[X;\mathcal{E}]$ to $C_{0}(\mathcal{G})$. In particular, $F^{p}[X;\mathcal{E}]$ is a groupoid Banach algebra. If $p<\infty$, then $\{1_{(y,j)}\}_{(y,j)\in X\times I}$ is a Schauder basis for $\ell^{p}(X,\mathcal{E})$, and then (7.14) implies that the map $j$ is injective on $F^{p}[X;\mathcal{E}]$ and hence $F^{P}[X;\mathcal{E}]$ is reduced. For $p=\infty$, this is not so clear. However, when $\mathcal{E}$ is reflexive, then $\{e_{i}^{*}\}_{i\in I}$ is a Schauder basis for the dual space $\mathcal{E}^{*}$, and the involution on $C_{c}(\mathcal{G})$ induces an isometric anti-linear isomorphism $F^{\infty}[X;\mathcal{E}]\cong F^{1}[X;\mathcal{E}^{*}]$. Thus $F^{\infty}[X;\mathcal{E}]$ is reduced because we have just proved that $F^{1}[X;\mathcal{E}^{*}]$ is. The assertion for the set $P\subseteq[1,\infty]$ follows now from Lemma 3.12.
@@
Using that $C_{0}(\beta(X)\times I)\subseteq F^{P}[X;\mathcal{E}]$ has the intersection property, one can show that ideals in $F^{P}[X;\mathcal{E}]$ in which controlled propagation operators are dense, are exactly those generated by open $\mathcal{G}$-invariant subsets of $\beta(X)\times I$ (which in turn correspond to ideals in the coarse structure $(X,\operatorname{\mathbb{E}})$). In this way, one could generalize [CWT04, Theorems 6.3, 6.4] from uniform Roe $C^{*}$-algebras to algebras of the form $F^{P}[X;\mathcal{E}]$.